Scores on a certain test were normally distributed with a mean of 80 and a standard deviation of ± 12. What is the probability that a given student got a score more than 80? Your answer should be correct to four decimal places.
Answer #1

Scores on a certain test ax ANN(80, 122) P(x)80) P ( X-80 > 80100) = P(azo) [X-MX = 2 ~N(0,1)] - 0.5ooo seare mare Probabilit